44 yd/s = mi/h
I got the answer .6 and it repeats. Just wondering if it's the right answer? If not explain why.

Answer by Alan3354(69443)

44 yd/s = mi/h
44 yd/s = 132 ft/sec
-----------
22 ft/sec = 15 mi/hr (you can confirm that)
132 = 6*22
--> 6*15 = 90 mi/hr